The information below pertains to Bonita Company for 2021.

Net income for the year $1,190,000
7% convertible bonds issued at par ($1,000 per bond); each bond is convertible into
    30 shares of common stock
1,970,000
6% convertible, cumulative preferred stock, $100 par value; each share is convertible
     into 3 shares of common stock
4,000,000
Common stock, $10 par value 6,280,000
Tax rate for 2021 20%
Average market price of common stock $25 per share


There were no changes during 2021 in the number of common shares, preferred shares, or convertible bonds outstanding. There is no treasury stock. The company also has common stock options (granted in a prior year) to purchase 68,800 shares of common stock at $20 per share.

(a) Compute basic earnings per share for 2021. (Round answer to 2 decimal places, e.g. $2.55.)

Basic earnings per share

$


(b) Compute diluted earnings per share for 2021. (Round answer to 2 decimal places, e.g. $2.55.)

Diluted earnings per share

$
Answer #1

Ans:

Net income = $1190000

Dividend on preference shares = 4000000*6%=

240000

Number of common stock outstanding = 6280000/10=628000

a Basic EPS = Net income-Preference Dividend/Common stock outstanding

=(1190000-240000)/628000= $1.51 (rounded off)

Net Income

1,190,000

DILUTED EPS

Interest on convertible bonds (1970,000*7%)

137900

After Tax amount(1-20%)

110320

Number of bonds

1970

Number of equivalent common shares @30 Anti dilution

59100

Stock Options

68800

Price

20

Value of Stock Options

1376000

Divided by Average Market price

25

Number of common stock

55040

Incremental shares

13760

Preference Shares

40000

Preference dividend

240,000

Number of Equivalent common shares@3

120000

Anti dilution

2

Numerator for Diluted EPS (Net Income- Preferred Dividend+Interest savings+ Preferred Dividend)

(1190,000+240000-240000+110320)

1,300,320

Total number of shares outstanding

820860

(628000+13760+59100+120000)

Diluted EPS

1.58